Determine the COP of a heat pump that supplies energy to a house at a rate of 8000 kJ/h for each kW of electric power it draws. Also, determine the rate of energy absorption from the outdoor air.

1
Expert's answer
2022-02-28T11:05:02-0500

Solution;

Given;

"\dot Q =8000kJ/h"

"W=1kW"

"COP_{HP}=\frac{Q}{W}=(\frac{8000kJ/h}{1kW})(\frac{1kW}{3600kJ/h})=2.22"

"\dot Q_L=\dot Q_{HP}-\dot W"

"\dot Q_L=8000-1(3600)=4400kJ/h"